OpenStudy (anonymous):

i need to complete the square of -36x^2-36x+27

OpenStudy (sasogeek):

have you tried solving this?

OpenStudy (anonymous):

yes. initially i divided each variable by 36, but then i got a bunch of fractions and it did not make sense to me. Do i need to divide by 36?

OpenStudy (sasogeek):

ok well that's a good start :) and yes you will end up with a bunch of fractions, it happens ;) i'll tell you what to do, try that, then let me know what answer you get... are we good?

OpenStudy (anonymous):

yes! please

OpenStudy (anonymous):

ok, i got x^2+x+(1/2)^2=-27/36-1/4

OpenStudy (sasogeek):

uhhh close, but not quite. you should've divided by -36 that way you'd have \(\large x^2+x=\frac{27}{36}\)

OpenStudy (sasogeek):

sorry for late reply, i lost connection

OpenStudy (anonymous):

ok, im going to try that. no problem, im doing some other problems while i wait

OpenStudy (sasogeek):

then with the next step, you'd get \(\large x^2+x+\frac{1}{4}=\frac{27}{36}+\frac{1}{4} \)

OpenStudy (anonymous):

ok, i have x^2+x+1/4=1/2

OpenStudy (sasogeek):

\(\large \frac{27}{36}+\frac{1}{4}= what? \)

OpenStudy (anonymous):

oops, i subtracted instead of adding. so 1?

OpenStudy (sasogeek):

yeah :)

OpenStudy (anonymous):

ok, now im a little confused, isnt it (x+b/2)^2. i dont know where to go from here

OpenStudy (sasogeek):

ok so you have \(\huge x^2+x+\frac{1}{4}=1 \) what you said is right, b=1, hence the next step is \(\huge (x+\frac{1}{2})^2=1\) :)

OpenStudy (anonymous):

so the answer is 1-(x+(1/2))^2

OpenStudy (sasogeek):

no... i keep losing connection so bear with the reply timing :) what you do is solve for x.... the next step is to find the square root of both sides of the equation to clear the square on the left side... \(\huge x+\frac{1}{2}= \pm \sqrt{1}\) \(\huge x= \pm \sqrt{1}-\frac{1}{2} \) \(\huge x= \pm 1-\frac{1}{2} \) solve for x now :)